# DreamGaussian: Generative Gaussian Splatting for Efficient 3D Content Creation

### v1

@@ 1 Introduction
On the other hand, Dreamfusion (Poole et al., 2022) proposes Score Distillation Sampling (SDS) to address the 3D data limitation by distilling 3D geometry and appearance from powerful 2D diffusion models (Saharia et al., 2022), which inspires the development of recent 2D lifting methods (Lin et al., 2023; Wang et al., 2023b; Chen et al., 2023b). In order to cope with the inconsistency and ambiguity caused by the SDS supervision, Neural Radiance Fields (NeRF) (Mildenhall et al., 2020) are usually adopted for their capability in modeling rich 3D information. Although the generation quality has been increasingly improved, these approaches are notorious for hours-long optimization time due to the costly NeRF rendering, which restricts them from being deployed to real-world applications at scale. We argue that the occupancy pruning technique used to accelerate NeRF (Müller et al., 2022; Sara Fridovich-Keil and Alex Yu et al., 2022) is ineffective in generative settings when supervised by the ambiguous SDS loss as opposed to reconstruction settings.
In this work, we introduce the DreamGaussian framework, which greatly improves the 3D content generation efficiency by refining the design choices in an optimization-based pipeline. Photo-realistic 3D assets with explicit mesh and texture maps can be generated from a single-view image within only 2 minutes using our method. Our core design is to adapt 3D Gaussian Splatting (Kerbl et al., 2023) into the generative setting with companioned meshes extraction and texture refinement. Compared to previous methods with the NeRF representation, which find difficulties in effectively pruning empty space, our generative Gaussian splatting significantly simplifies the optimization landscape. Specifically, we demonstrate the progressive densification of Gaussian splitting, which highly in accordance with the optimization progress of generative settings, greatly improves the generation efficiency. As illustrated in Figure 1, our image-to-3D pipeline swiftly produces a coarse shape within seconds and converges efficiently in around $500$ steps on a single GPU.
@@
    We propose a novel framework for 3D content creation by adapting Gaussian splatting into generative settings, significantly reducing the generation time of optimization-based 2D lifting methods.
@@ 2 Related Work
Various 3D representations have been proposed for different 3D tasks. Neural Radiance Fields (NeRF) (Mildenhall et al., 2020) employs a volumetric rendering and has been popular for enabling 3D optimization with only 2D supervision. Although NeRF has become widely used in both 3D reconstruction (Barron et al., 2022; Li et al., 2023c; Chen et al., 2022; Hedman et al., 2021) and generation (Poole et al., 2022; Lin et al., 2023; Chan et al., 2022), optimizing NeRF can be time-consuming. Various attempts have been made to accelerate the training of NeRF (Müller et al., 2022; Sara Fridovich-Keil and Alex Yu et al., 2022), but these works only focus on the reconstruction setting. The common technique of spatial pruning fails to accelerate the generation setting. Recently, 3D Gaussian splatting (Kerbl et al., 2023) has been proposed as an alternative 3D representation to NeRF, which has demonstrated impressive quality and speed in 3D reconstruction (Luiten et al., 2023). The efficient differentiable rendering implementation and model design enables fast training without relying on spatial pruning. In this work, we for the first time adapt 3D Gaussian splatting into generation tasks to release the potential of optimization-based methods.
@@ 2.2 Text-to-3D Generation
Text-to-3D generation aims at generating 3D assets from a text prompt. Recently, data-driven 2D diffusion models have achieved notable success in text-to-image generation (Ho et al., 2020; Rombach et al., 2022; Saharia et al., 2022). However, transferring it to 3D generation is non-trivial due to the challenge of curating large-scale 3D datasets. Existing 3D native diffusion models usually work on a single object category and suffer from limited diversity (Jun & Nichol, 2023; Nichol et al., 2022; Gupta et al., 2023; Lorraine et al., 2023; Zhang et al., 2023; Zheng et al., 2023; Ntavelis et al., 2023; Chen et al., 2023a; Cheng et al., 2023; Gao et al., 2022). To achieve open-vocabulary 3D generation, severl methods propose to lift 2D image models for 3D generation (Jain et al., 2022; Poole et al., 2022; Wang et al., 2023a; Mohammad Khalid et al., 2022; Michel et al., 2022). Such 2D lifting methods optimize a 3D representation to achieve a high likelihood in pretrained 2D diffusion models when rendered from different viewpoints, such that both 3D consistency and realisticity can be ensured. Following works continue to enhance various aspects such as generation fidelity and training stability (Lin et al., 2023; Tsalicoglou et al., 2023; Zhu & Zhuang, 2023; Yu et al., 2023; Li et al., 2023b; Chen et al., 2023c; Wang et al., 2023b; Huang et al., 2023; Metzer et al., 2022; Chen et al., 2023b), and explore further applications (Zhuang et al., 2023; Singer et al., 2023; Raj et al., 2023). However, these optimization-based 2D lifting approaches usually suffer from long per-case optimization time. Particularly, employing NeRF as the 3D representation leads to expensive computations during both forward and backward. In this work, we choose 3D Gaussians as the differentiable 3D representation and empirically show that it has a simpler optimization landscape.
@@ 3 Our Approach
In this section, we introduce our two-stage framework for efficient 3D content generation for both Image-to-3D and Text-to-3D tasks as illustrated in Figure 2. Firstly, we adapt 3D Gaussian splatting (Kerbl et al., 2023) into generation tasks for efficient initialization through SDS (Poole et al., 2022) (Section 3.1). Next, we propose an efficient algorithm to extract a textured mesh from 3D Gaussians (Section 3.2). This texture is then fine-tuned by differentiable rendering (Laine et al., 2020) through a UV-space refinement stage (Section 3.3) for final exportation.
@@ 3.1 Generative Gaussian Splatting
Gaussian splatting (Kerbl et al., 2023) represents 3D information with a set of 3D Gaussians. It has been proven effective in reconstruction settings (Kerbl et al., 2023; Luiten et al., 2023) with high inference speed and reconstruction quality under similar modeling time with NeRF. However, its usage in a generative manner has not been explored. We identify that the 3D Gaussians can be efficient for 3D generation tasks.
Specifically, the location of each Gaussian can be described with a center $\mathbf{x}\in\mathbb{R}^{3}$, a scaling factor $\mathbf{s}\in\mathbb{R}^{3}$, and a rotation quaternion $\mathbf{q}\in\mathbb{R}^{4}$. We also store an opacity value $\alpha\in\mathbb{R}$ and a color feature $\mathbf{c}\in\mathbb{R}^{3}$ for volumetric rendering. Spherical harmonics are disabled since we only want to model simple diffuse color. All the above optimizable parameters is presented by ${\Theta}$, where ${\Theta}_{i}=\{\mathbf{x}_{i},\mathbf{s}_{i},\mathbf{q}_{i},\alpha_{i},\mathbf{c}_{i}\}$ is the parameter for the $i$-th Gaussian. To render a set of 3D Gaussians, we need to project them into the image plane as 2D Gaussians. Volumetric rendering is then performed for each pixel in front-to-back depth order to evaluate the final color and alpha. In this work, we use the highly optimized renderer implementation from Kerbl et al. (2023) to optimize ${\Theta}$.
We initialize the 3D Gaussians with random positions sampled inside a sphere, with unit scaling and no rotation. These 3D Gaussians are periodically densified during optimization. Different from the reconstruction pipeline, we start from fewer Gaussians but densify it more frequently to align with the generation progress. We follow the recommended practices from previous works (Poole et al., 2022; Huang et al., 2023; Lin et al., 2023) and use SDS to optimize the 3D Gaussians. At each step, we sample a random camera pose $p$ orbiting the object center, and render the RGB image $I^{p}_{\text{RGB}}$ and transparency $I^{p}_{\text{A}}$ of the current view. Similar to Dreamtime (Huang et al., 2023), we decrease the timestep $t$ linearly, which is used to weight the random noise $\epsilon$ added to the rendered RGB image. Then, different 2D diffusion priors $\phi$ are used to guide the SDS denoising step, which is back-propagated to the 3D Gaussians.
@@
$\nabla_{\Theta}\mathcal{L}_{\text{SDS}}=\mathbb{E}_{t,p,\mathbf{\epsilon}}\left[(\epsilon_{\phi}(I^{p}_{\text{RGB}};t,\tilde{I}^{r}_{\text{RGB}},\Delta p)-\epsilon)\frac{\partial I^{p}_{\text{RGB}}}{\partial{\Theta}}\right]$ (1)
where $\epsilon_{\phi}(\cdot)$ is the predicted noise by the 2D diffusion prior $\phi$, and $\Delta p$ is the relative camera pose change from the reference camera $r$. Additionally, we optimize the reference view image $I_{\text{RGB}}^{r}$ and transparency $I_{\text{A}}^{r}$ to align with the input:
@@
$\nabla_{\Theta}\mathcal{L}_{\text{SDS}}=\mathbb{E}_{t,p,\mathbf{\epsilon}}\left[(\epsilon_{\phi}(I^{p}_{\text{RGB}};t,e)-\epsilon)\frac{\partial I^{p}_{\text{RGB}}}{\partial{\Theta}}\right]$ (3)
@@
Dissusion. We observe that the generated Gaussians often look blurry and lack details even with longer SDS training iterations. This could be explained by the ambiguity of SDS loss. Since each optimization step may provide inconsistent 3D guidance, it’s hard for the algorithm to correctly densify the under-reconstruction regions or prune over-reconstruction regions as in reconstruction. This observation leads us to the following mesh extraction and texture refinement designs.
@@ 3.2 Efficient Mesh Extraction
We first divide the 3D space of $(-1,1)^{3}$ into $16^{3}$ blocks, then cull the Gaussians whose centers are located outside each local block. This effectively reduces the total number of Gaussians to query in each block. We then query a $8^{3}$ dense grid inside each block, which leads to a final $128^{3}$ dense grid. For each query at grid position $\mathbf{x}$, we sum up the weighted opacity of each remained 3D Gaussian:
@@
where $\Sigma_{i}$ is the covariance matrix built from scaling $\mathbf{s}_{i}$ and rotation $\mathbf{q}_{i}$. An empirical threshold is then used to extract the mesh surface through Marching Cubes. Decimation and Remeshing (Cignoni et al., 2008) are applied to post-process the extracted mesh to make it smooth.
Color Back-projection. Since we have acquired the mesh geometry, we can back project the rendered RGB image to the mesh surface and bake it as the texture. We first unwrap the mesh’s UV coordinates (Young, 2021) and initialize an empty texture image. Then we uniformly choose 8 azimuths and 3 elevations, plus the top and bottom views to render the corresponding RGB image. Each pixel from these RGB images can be back-projected to the texture image based on the UV coordinate. Following Richardson et al. (2023), we exclude the pixels with a small camera space z-direction normal to avoid unstable projection at mesh boundaries. This back-projected texture image serves as an initialization for the next mesh texture fine-tuning stage.
@@ 3.3 UV-space Texture Refinement
Due to the ambiguity of SDS optimization, the meshes extracted from 3D Gaussians usually possess blurry texture as illustrated in Figure 3. We therefore propose a second stage to refine the texture image. However, fine-tuning the UV-space directly with SDS loss often leads to artifacts, which is also observed in previous works (Liao et al., 2023). This is due to the mipmap texture sampling technique used in differentiable rasterization (Laine et al., 2020). With ambiguous guidance like SDS, the gradient propagated to each mipmap level results in over-saturated color blocks. Therefore, we seek more definite guidance to fine-tune a blurry texture.
@@ 4.1 Implementation Details
We train $500$ steps for the first stage and $50$ steps for the second stage. The 3D Gaussians are initialized to $0.1$ opacity and grey color inside a sphere of radius $0.5$. The rendering resolution is increased from $64$ to $512$ for Gaussian splatting, and randomly sampled from $128$ to $1024$ for mesh. The loss weights for RGB and transperency are linearly increased from 00 to $10^{4}$ and $10^{3}$ during training. We sample random camera poses at a fixed radius of $2$ for image-to-3D and $2.5$ for text-to-3D, y-axis FOV of $49$ degree, with the azimuth in $[-180,180]$ degree and elevation in $[-30,30]$ degree. The background is rendered randomly as white or black for Gaussian splatting. For image-to-3D task, the two stages each take around 1 minute. We preprocess the input image by background removal (Qin et al., 2020) and recenter the foreground object. The 3D Gaussians are initialized with $5000$ random particles and densified for each $100$ steps. For text-to-3D task, due to the larger resolution of $512\times 512$ used by Stable Diffusion (Rombach et al., 2022) model, each stage takes around 2 minutes to finish. We initialize the 3D Gaussians with $1000$ random particles and densify them for each $50$ steps. For mesh extraction, we use an empirical threshold of $1$ for Marching Cubes. All experiments are performed and measured with an NVIDIA V100 (16GB) GPU, while our method requires less than 8 GB GPU memory. Please check the supplementary materials for more details.
@@
$*{}^{*}$ (Liu et al., 2023b) Optimization-based 0.778 30 minutes
@@
Table 1: Quantitative Comparisons on generation quality and speed for image-to-3D tasks. For $*{}^{*}$, a mesh fine-tuning stage is used to further improve quality (Tang, 2022).

### v2

@@ 1 Introduction
On the other hand, Dreamfusion (Poole et al., 2022) proposes Score Distillation Sampling (SDS) to address the 3D data limitation by distilling 3D geometry and appearance from powerful 2D diffusion models (Saharia et al., 2022), which inspires the development of recent 2D lifting methods (Lin et al., 2023; Wang et al., 2023b; Chen et al., 2023c). In order to cope with the inconsistency and ambiguity caused by the SDS supervision, Neural Radiance Fields (NeRF) (Mildenhall et al., 2020) are usually adopted for their capability in modeling rich 3D information. Although the generation quality has been increasingly improved, these approaches are notorious for hours-long optimization time due to the costly NeRF rendering, which restricts them from being deployed to real-world applications at scale. We argue that the occupancy pruning technique used to accelerate NeRF (Müller et al., 2022; Sara Fridovich-Keil and Alex Yu et al., 2022) is ineffective in generative settings when supervised by the ambiguous SDS loss as opposed to reconstruction settings.
In this work, we introduce the DreamGaussian framework, which greatly improves the 3D content generation efficiency by refining the design choices in an optimization-based pipeline. Photo-realistic 3D assets with explicit mesh and texture maps can be generated from a single-view image within only 2 minutes using our method. Our core design is to adapt 3D Gaussian Splatting (Kerbl et al., 2023) into the generative setting with companioned meshes extraction and texture refinement. Compared to previous methods with the NeRF representation, which find difficulties in effectively pruning empty space, our generative Gaussian splatting significantly simplifies the optimization landscape. Specifically, we demonstrate the progressive densification of Gaussian splatting, which is in accordance with the optimization progress of generative settings, greatly improves the generation efficiency. As illustrated in Figure 1, our image-to-3D pipeline swiftly produces a coarse shape within seconds and converges efficiently in around $500$ steps on a single GPU.
@@
    We adapt 3D Gaussian splatting into generative settings for 3D content creation, significantly reducing the generation time of optimization-based 2D lifting methods.
@@ 2 Related Work
Various 3D representations have been proposed for different 3D tasks. Neural Radiance Fields (NeRF) (Mildenhall et al., 2020) employs a volumetric rendering and has been popular for enabling 3D optimization with only 2D supervision. Although NeRF has become widely used in both 3D reconstruction (Barron et al., 2022; Li et al., 2023d; Chen et al., 2022; Hedman et al., 2021) and generation (Poole et al., 2022; Lin et al., 2023; Chan et al., 2022), optimizing NeRF can be time-consuming. Various attempts have been made to accelerate the training of NeRF (Müller et al., 2022; Sara Fridovich-Keil and Alex Yu et al., 2022), but these works mostly focus on the reconstruction setting. The common technique of spatial pruning fails to accelerate the generation setting. Recently, 3D Gaussian splatting (Kerbl et al., 2023) has been proposed as an alternative 3D representation to NeRF, which has demonstrated impressive quality and speed in 3D reconstruction (Luiten et al., 2023). The efficient differentiable rendering implementation and model design enables fast training without relying on spatial pruning. In this work, we for the first time adapt 3D Gaussian splatting into generation tasks to unlock the potential of optimization-based methods.
@@ 2.2 Text-to-3D Generation
Text-to-3D generation aims at generating 3D assets from a text prompt. Recently, data-driven 2D diffusion models have achieved notable success in text-to-image generation (Ho et al., 2020; Rombach et al., 2022; Saharia et al., 2022). However, transferring it to 3D generation is non-trivial due to the challenge of curating large-scale 3D datasets. Existing 3D native diffusion models usually work on a single object category and suffer from limited diversity (Jun & Nichol, 2023; Nichol et al., 2022; Gupta et al., 2023; Lorraine et al., 2023; Zhang et al., 2023; Zheng et al., 2023; Ntavelis et al., 2023; Chen et al., 2023b; Cheng et al., 2023; Gao et al., 2022). To achieve open-vocabulary 3D generation, several methods propose to lift 2D image models for 3D generation (Jain et al., 2022; Poole et al., 2022; Wang et al., 2023a; Mohammad Khalid et al., 2022; Michel et al., 2022). Such 2D lifting methods optimize a 3D representation to achieve a high likelihood in pretrained 2D diffusion models when rendered from different viewpoints, such that both 3D consistency and realisticity can be ensured. Following works continue to enhance various aspects such as generation fidelity and training stability (Lin et al., 2023; Tsalicoglou et al., 2023; Zhu & Zhuang, 2023; Yu et al., 2023; Li et al., 2023c; Chen et al., 2023d; Wang et al., 2023b; Huang et al., 2023; Metzer et al., 2022; Chen et al., 2023c), and explore further applications (Zhuang et al., 2023; Singer et al., 2023; Raj et al., 2023). However, these optimization-based 2D lifting approaches usually suffer from long per-case optimization time. Particularly, employing NeRF as the 3D representation leads to expensive computations during both forward and backward. In this work, we choose 3D Gaussians as the differentiable 3D representation and empirically show that it has a simpler optimization landscape.
@@ 3 Our Approach
In this section, we introduce our two-stage framework for efficient 3D content generation for both Image-to-3D and Text-to-3D tasks as illustrated in Figure 2. Firstly, we adapt 3D Gaussian splatting (Kerbl et al., 2023) into generation tasks for efficient initialization through SDS (Poole et al., 2022) (Section 3.1). Next, we propose an algorithm to extract a textured mesh from 3D Gaussians (Section 3.2). This texture is then fine-tuned by differentiable rendering (Laine et al., 2020) through a UV-space refinement stage (Section 3.3) for final exportation.
@@ 3.1 Generative Gaussian Splatting
Gaussian splatting (Kerbl et al., 2023) represents 3D information with a set of 3D Gaussians. It has been proven effective in reconstruction settings (Kerbl et al., 2023; Luiten et al., 2023) with high inference speed and reconstruction quality under similar modeling time with NeRF. However, its usage in a generative manner has not been explored. We identify that the 3D Gaussians can be efficient for 3D generation tasks too.
Specifically, the location of each Gaussian can be described with a center $\mathbf{x}\in\mathbb{R}^{3}$, a scaling factor $\mathbf{s}\in\mathbb{R}^{3}$, and a rotation quaternion $\mathbf{q}\in\mathbb{R}^{4}$. We also store an opacity value $\alpha\in\mathbb{R}$ and a color feature $\mathbf{c}\in\mathbb{R}^{3}$ for volumetric rendering. Spherical harmonics are disabled since we only want to model simple diffuse color. All the above optimizable parameters is presented by ${\Theta}$, where ${\Theta}_{i}=\{\mathbf{x}_{i},\mathbf{s}_{i},\mathbf{q}_{i},\alpha_{i},\mathbf{c}_{i}\}$ is the parameter for the $i$-th Gaussian. To render a set of 3D Gaussians, we need to project them onto the image plane as 2D Gaussians. Volumetric rendering is then performed for each pixel in front-to-back depth order to evaluate the final color and alpha. In this work, we use the highly optimized renderer implementation from Kerbl et al. (2023) to optimize ${\Theta}$.
We initialize the 3D Gaussians with random positions sampled inside a sphere, with unit scaling and no rotation. These 3D Gaussians are periodically densified during optimization. Different from the reconstruction pipeline, we start from fewer Gaussians but densify it more frequently to align with the generation progress. We follow the recommended practices from previous works (Poole et al., 2022; Huang et al., 2023; Lin et al., 2023) and use SDS to optimize the 3D Gaussians (Please refer to Section A.1 for more details on SDS loss). At each step, we sample a random camera pose $p$ orbiting the object center, and render the RGB image $I^{p}_{\text{RGB}}$ and transparency $I^{p}_{\text{A}}$ of the current view. Similar to Dreamtime (Huang et al., 2023), we decrease the timestep $t$ linearly during training, which is used to weight the random noise $\epsilon$ added to the rendered RGB image. Then, different 2D diffusion priors $\phi$ can be used to optimize the underlying 3D Gaussians through SDS.
@@
$\nabla_{\Theta}\mathcal{L}_{\text{SDS}}=\mathbb{E}_{t,p,\mathbf{\epsilon}}\left[w(t)(\epsilon_{\phi}(I^{p}_{\text{RGB}};t,\tilde{I}^{r}_{\text{RGB}},\Delta p)-\epsilon)\frac{\partial I^{p}_{\text{RGB}}}{\partial{\Theta}}\right]$ (1)
where $w(t)$ is a weighting function, $\epsilon_{\phi}(\cdot)$ is the predicted noise by the 2D diffusion prior $\phi$, and $\Delta p$ is the relative camera pose change from the reference camera $r$. Additionally, we optimize the reference view image $I_{\text{RGB}}^{r}$ and transparency $I_{\text{A}}^{r}$ to align with the input:
@@
$\nabla_{\Theta}\mathcal{L}_{\text{SDS}}=\mathbb{E}_{t,p,\mathbf{\epsilon}}\left[w(t)(\epsilon_{\phi}(I^{p}_{\text{RGB}};t,e)-\epsilon)\frac{\partial I^{p}_{\text{RGB}}}{\partial{\Theta}}\right]$ (3)
@@
Discussion. We observe that the generated Gaussians often look blurry and lack details even with longer SDS training iterations. This could be explained by the ambiguity of SDS loss. Since each optimization step may provide inconsistent 3D guidance, it’s hard for the algorithm to correctly densify the under-reconstruction regions or prune over-reconstruction regions as in reconstruction. This observation leads us to the following mesh extraction and texture refinement designs.
@@ 3.2 Efficient Mesh Extraction
We first divide the 3D space of $(-1,1)^{3}$ into $16^{3}$ overlapping blocks, then cull the Gaussians whose centers are located outside each local block. This effectively reduces the total number of Gaussians to query in each block. We then query a $8^{3}$ dense grid inside each block, which leads to a final $128^{3}$ dense grid. For each query at grid position $\mathbf{x}$, we sum up the weighted opacity of each remained 3D Gaussian:
@@
where $\Sigma_{i}$ is the covariance matrix built from scaling $\mathbf{s}_{i}$ and rotation $\mathbf{q}_{i}$. An empirical threshold is then used to extract the mesh surface through Marching Cubes. Decimation and remeshing (Cignoni et al., 2008) are applied to post-process the extracted mesh to make it smoother and more compact.
Color Back-projection. Since we have acquired the mesh geometry, we can back-project the rendered RGB image to the mesh surface and bake it as the texture. We first unwrap the mesh’s UV coordinates (Young, 2021) (detailed in Section A.1) and initialize an empty texture image. Then, we uniformly choose 8 azimuths and 3 elevations, plus the top and bottom views to render the corresponding RGB image. Each pixel from these RGB images can be back-projected to the texture image based on its UV coordinate. Following Richardson et al. (2023), we exclude the pixels with a small camera space z-direction normal to avoid unstable projection at mesh boundaries. This back-projected texture image serves as an initialization for the next texture fine-tuning stage.
@@ 3.3 UV-space Texture Refinement
We further use a second stage to refine the extracted coarse texture. Different from texture generation (Richardson et al., 2023; Chen et al., 2023a; Cao et al., 2023), we hope to enhance the details given a coarse texture. However, fine-tuning the UV-space directly with SDS loss leads to artifacts as shown in Figure 3, which is also observed in previous works (Liao et al., 2023). This is due to the mipmap texture sampling technique used in differentiable rasterization (Laine et al., 2020). With ambiguous guidance like SDS, the gradient propagated to each mipmap level results in over-saturated color blocks. Therefore, we seek more definite guidance to fine-tune a blurry texture.
@@ 4.1 Implementation Details
We train $500$ steps for the first stage and $50$ steps for the second stage. The 3D Gaussians are initialized to $0.1$ opacity and grey color inside a sphere of radius $0.5$. The rendering resolution is increased from $64$ to $512$ for Gaussian splatting, and randomly sampled from $128$ to $1024$ for mesh. The loss weights for RGB and transperency are linearly increased from 00 to $10^{4}$ and $10^{3}$ during training. We sample random camera poses at a fixed radius of $2$ for image-to-3D and $2.5$ for text-to-3D, y-axis FOV of $49$ degree, with the azimuth in $[-180,180]$ degree and elevation in $[-30,30]$ degree. The background is rendered randomly as white or black for Gaussian splatting. For image-to-3D task, the two stages each take around 1 minute. We preprocess the input image by background removal (Qin et al., 2020) and recentering of the foreground object. The 3D Gaussians are initialized with $5000$ random particles and densified for each $100$ steps. For text-to-3D task, due to the larger resolution of $512\times 512$ used by Stable Diffusion (Rombach et al., 2022) model, each stage takes around 2 minutes to finish. We initialize the 3D Gaussians with $1000$ random particles and densify them for each $50$ steps. For mesh extraction, we use an empirical threshold of $1$ for Marching Cubes. All experiments are performed and measured with an NVIDIA V100 (16GB) GPU, while our method requires less than 8 GB GPU memory. Please check the supplementary materials for more details.
@@
Zero-1-to-3 (Liu et al., 2023b) Optimization-based 0.778 30 minutes
@@
Table 1: Quantitative Comparisons on generation quality and speed for image-to-3D tasks. For Zero-1-to-3, a mesh fine-tuning stage is used to further improve quality (Tang, 2022).